Write the coordinates for the center of the hyperbola:

((y - 2) ^ 2)/16 - ((x + 1) ^ 2)/144 = 1


1
Expert's answer
2021-10-06T16:51:30-0400

Solution.

Equation of hyperbola in general form:


"\frac{(x-m)^{2}}{a^{2}}-\frac{(y-n)^{2}}{b^{2}}=1"


where

m, n – coordinates of the center of the hyperbola

"a, b \in \mathbb {R}"


The given hyperbola in the problem is rotated to "\frac{\pi}{2}":

"\frac{(y-2)^{2}}{4^{2}}-\frac{(x+1)^{2}}{12^{2}}=1"



Figure 1. Hyperbola rotated by "\frac{\pi}{2}"

"\begin{matrix}
 m = -1 \\
 n = 2
\end{matrix} \implies" Hyperbola with center (-1, 2)


Answer: (-1, 2)